# Cavity Control of Topological Qubits: Fusion Rule, Anyon Braiding and Majorana-Schrödinger Cat States

### v1

@@ Abstract
Majorana-Schrödinger cat states.— Photon Schrödinger cat states have been intensively studied in circuit quantum electrodynamics scenarios since they offer the opportunity of investigating fundamental tests of quantum theory, for encoding logical qubits and for developing error correction codes [42]. A Schrödinger cat-like state can be written as $\lvert{\rm\bf Cat}_{\pm}\rangle=\left(\lvert\alpha\rangle\pm\lvert-\alpha\rangle\right)/N_{\pm}$ with normalization constant $N_{\pm}^{2}=2\left(1\pm\exp\left[-2|\alpha|^{2}\right]\right)$ and sign $+$ (--) for even (odd) cat state. Interestingly enough, in the USC regime, a $2CK$ system supports photon Schrödinger cat states coupled to Majorana fermion states at the special coupling values $\lambda_{1}=\lambda_{2}=\lambda$, $\phi_{1}=0$, and $\phi_{2}=\pi$. The $2CK$ ground state becomes:
$\begin{split}\lvert{\rm GS}\rangle&\simeq\lvert+,-\rangle_{x}\otimes|-\alpha\rangle+\lvert-,+\rangle_{x}\otimes|\alpha\rangle,\\ &\simeq\lvert\tilde{\Psi}_{+}\rangle\otimes\lvert{\rm\bf Cat}_{+}\rangle+\lvert\tilde{\Psi}_{-}\rangle\otimes\lvert{\rm\bf Cat}_{-}\rangle.\end{split}$ (6)
@@
Where, we are using $\alpha=-\lambda/(\sqrt{2}\omega)$, $\lvert\tilde{\Psi}_{\pm}\rangle=(\lvert\bullet_{L,R},\circ_{c_{1}},\bullet_{c_{2}}\rangle\pm\lvert\bullet_{L,R},\bullet_{c_{1}},\circ_{c_{2}}\rangle)/\sqrt{2}$, and the matter basis set $\{|n_{L,R},n_{c_{1}},n_{c_{2}}\rangle\}$ has been used (number occupation representation of the non-local fermion $f_{L,R}$ and two local fermions fully contained inside the cavity, $c_{1}$ and $c_{2}$, see SM [28]). The result in Eq. (6) indicates that the system’s ground state is an entangled KC-cavity state: the symmetric combination of the two local fermion ($c_{1},c_{2}$) state couples with an even Schrödinger cat state whereas the anti-symmetric two fermion state couples to the odd Schrödinger cat state. Therefore, conditioned on the symmetry measurement outcome of the fermion state inside the $2CK$, an even or odd cavity cat state can be distilled. In the resonant case, the ground state of the Hamiltonian given by Eq. (2) is numerically obtained through exact diagonalization from which the reduced photon density matrix immediately is within reach for the GS out of the special point $\lambda_{1}=\lambda_{2}=\lambda$, $\phi_{1}=0$, and $\phi_{2}=\pi$. As depicted in the upper-lateral plots in Fig. 5 the Wigner functions do not display negative features as expected for coherent states. However, for the special set of coupling strengths we are considering in Eq. (6), after a previous projection of the GS on the matter states $\lvert\tilde{\Psi}_{\pm}\rangle$, the corresponding Wigner function displays negative values in the central plots of Fig. 5, a fingerprint of Majorana induced Schrödinger cat states.
@@
Figure 5: Majorana-Schrödinger cat states. In the main panels (a) and (b), the bosonic Wigner function is represented, conditioned on the symmetry measurement outcome of the fermion state inside the 2CK in the resonance case. An even (panel (a)) or odd (panel (b)) cavity cat state is distilled. The red lines over the density projections correspond to the values of $\alpha=\pm\lambda/(\sqrt{2}\omega)$, with $\lambda=4$ fixed. Similarly, the insets correspond to the numerical Wigner functions obtained when the ground state is projected onto $\left|+,-\right\rangle_{x}$ and $\left|-,+\right\rangle_{x}$, respectively.
Summary and discussion.— In this work, we report on a diverse range of new features of anyons that can be significant for harnessing cavity confinement control over quantum topological materials. Our results not only uncover new facets of anyon-photon composed states at a fundamental level but also have interesting applications, such as braiding-allowed teleportation schemes in the USC regime. Our findings provide useful insights into how superconducting circuit cavity control is indeed a feasible alternative to test non-Abelian anyon statistics associated with MZMs. This would provide a stringent test of the theory of free Majorana fermions in nanowire or quantum dot array scenarios.
Enhanced experimental capabilities to control superconducting quantum electrodynamics circuits are currently fuelling intense research in the domain of light-matter systems. Harnessing ultra-strong couplings is thus not only of fundamental interest but also of appeal for technological applications. Regarding the experimental feasibility of the above-discussed results, it is worth noting that the KC model has been demonstrated experimentally in a fully controlled chain of coupled quantum dots where fine-tuning of parameters lays out the system into a sweet spot which leads to the generation of poor man’s Majorana states [6]. The next step should be the coupling to a transmission line resonator. With properly applied driving on the resonator mode or directly on the KC sites embedded within the cavity, the cavity-controlled KC features can be observed from the transmission spectrum [23, 24, 36, 37], by quantum process tomography [38] or by microwave interferometric measurements [39].
An interesting prospect is to generalize these single cavity controlled anyon processes to multi-cavity setups offering tools to observe geometric phase controlled braiding mechanisms offering new paths for further exploration. Besides, dissipation is something that cannot be avoided in experiments and thus should be accounted for in studying the photon-matter correlation transitions.
@@
Acknowledgments.— L.Q. and F.J.R. would like to thank Facultad de Ciencias-UniAndes projects: INV-2021-128-2292 and INV-2023-162-2833 for financial support. F.J.G-R and I.A.B-G acknowledge financial support from Spanish MCIN with funding from European Union Next Generation EU (PRTRC17.I1) and Consejeria de Educacion from Junta de Castilla y Leon through QCAYLE project, as well as grants PID2020-113406GB-I00 MTM funded by AEI/10.13039/501100011033, and RED2022-134301-T. C.T. acknowledges financial support from the project “Nanofotonica para Computacion Cuantica” (NanoQuCo) with reference Y2020/TCS-6545 of the Comunidad de Madrid (CAM).

### v2

@@ 5 Majorana-Schrödinger cat states
Photon Schrödinger cat states have been intensively studied in circuit quantum electrodynamics scenarios since they offer the opportunity of investigating fundamental tests of quantum theory, for encoding logical qubits and for developing error correction codes [65]. A Schrödinger cat-like state can be written as $|{\rm\bf Cat}_{\pm}\rangle=\left(|\alpha\rangle\pm|-\alpha\rangle\right)/N_{\pm}$ with normalization constant $N_{\pm}^{2}=2\left(1\pm\exp\left[-2|\alpha|^{2}\right]\right)$ and sign ++ (-) for even (odd) cat state. Interestingly enough, in the USC regime, a $2CK$ system supports photon Schrödinger cat states coupled to Majorana fermion states at the special coupling values $\lambda_{1}=\lambda_{2}=\lambda$, $\phi_{1}=0$, and $\phi_{2}=\pi$. The $2CK$ ground state becomes:
$|{\rm GS}\rangle$ $\simeq|+,-\rangle_{x}\otimes|-\alpha\rangle+|-,+\rangle_{x}\otimes|\alpha\rangle,$ (14)
$\simeq|\tilde{\Psi}_{+}\rangle\otimes|{\rm\bf Cat}_{+}\rangle+|\tilde{\Psi}_{-}\rangle\otimes|{\rm\bf Cat}_{-}\rangle.$ (15)
@@
Where, we are using $\alpha=-\lambda/(\sqrt{2}\omega)$, $|\tilde{\Psi}_{\pm}\rangle=(|\bullet_{L,R},\circ_{c_{1}},\bullet_{c_{2}}\rangle\pm|\bullet_{L,R},\bullet_{c_{1}},\circ_{c_{2}}\rangle)/\sqrt{2}$, and the matter basis set $\{|n_{L,R},n_{c_{1}},n_{c_{2}}\rangle\}$ has been used (number occupation representation of the non-local fermion $f_{L,R}$ and two local fermions fully contained inside the cavity, $c_{1}$ and $c_{2}$, see 2.2). The result in Eq. (14) indicates that the system’s ground state is an entangled KC-cavity state: the symmetric combination of the two local fermion ($c_{1},c_{2}$) state couples with an even Schrödinger cat state whereas the anti-symmetric two fermion state couples to the odd Schrödinger cat state. Therefore, conditioned on the symmetry measurement outcome of the fermion state inside the $2CK$, an even or odd cavity cat state can be distilled. In the resonant case, the ground state of the Hamiltonian given by Eq. (9) is numerically obtained through exact diagonalization from which the reduced photon density matrix immediately is within reach for the GS out of the special point $\lambda_{1}=\lambda_{2}=\lambda$, $\phi_{1}=0$, and $\phi_{2}=\pi$. As depicted in the upper-lateral plots in Fig. 5 the Wigner functions do not display negative features as expected for coherent states. However, for the special set of coupling strengths we are considering in Eq. (14), after a previous projection of the GS on the matter states $|\tilde{\Psi}_{\pm}\rangle$, the corresponding Wigner function displays negative values in the central plots of Fig. 5, a fingerprint of Majorana induced Schrödinger cat states.
The photon cat states arise specifically due to the cavity scissor effect in a $2CK$ system. This effect enables the formation of free, self-adjoint Majorana zero modes within the KC bulk, enabled by the cavity-induced splitting of the chain. These MZMs are responsible for encoding non-local topological qubits, which, in turn, induce the observed photon cat states. This unique interplay between the KC and the cavity could offer a reliable mean to indirectly detect the presence of MZMs through optical signatures. Moreover, mapping the full Hamiltonian into an effective, yet exact, spin-photon model, we dispose of an efficient and exact method to simplify the problem without sacrificing its key physical insights. This Rabi-like mapping has proven to be a powerful tool for extracting non-trivial results, such as the photon cat states, in a clear and computationally manageable framework.
@@
[…]
6 Summary and discussion
In this work, we addressed a conceptual question: can a single- or double-site cavity embedded in a Kitaev chain effectively extract and control the chain’s topological features? By analyzing this setup, we uncovered a rich interplay between cavity quantum electrodynamics and topological quantum matter. We demonstrated that the cavity acts not merely as a passive boundary splitting the chain, but as an active quantum component with its own dynamical degrees of freedom. This makes it essential for generating and mediating non-local quantum correlations between photons and matter.
Our results show that embedding a KC within a spatially selective cavity in the ultrastrong coupling regime enables the realization of non-Abelian anyon properties, such as fusion rules and braiding operations, within a fully one-dimensional platform. These phenomena can be probed through experimentally viable signatures like fermionic parity readouts and photon-induced Berry phases. Additionally, we showed that hybrid Majorana-polariton Schrödinger cat states can be generated via the symmetry of fermionic modes interacting within a two-site cavity configuration, opening a novel path for topological quantum information processing.
Importantly, we demonstrated that in the so-called sweet spot of the Kitaev chain, often realized in quantum dot arrays, the coupled light-matter Hamiltonian reduces to a Rabi-like model with rotated homodyne quadratures. This mapping simplifies the theoretical treatment and provides a direct link to experimental observables in cavity QED architectures. Furthermore, deviations from the ideal $\mu=0$ condition do not critically affect the non-local character of the Majorana modes, as supported by prior studies [25], indicating that fine-tuning requirements can be relaxed.
While our setup exhibits compelling features and supports various non-Abelian signatures, we also acknowledge some of its limitations. In particular, the lack of topological protection in the minimal Kitaev chains ($1CK$ and $2CK$) implies sensitivity to disorder and decoherence, which should be carefully addressed in future experimental implementations. Nevertheless, the ability to control the cavity parameters and the accessibility of measurable observables still make this platform an attractive and realistic testbed for probing Majorana physics.
We also emphasize that our focus has been on proposing a proof-of-principle platform for photon-assisted anyonic phenomena, rather than presenting a fault-tolerant topological quantum computing architecture. As such, our findings should be viewed as a step toward merging cavity QED and topological phases of matter, with potential to evolve into more robust schemes.
Our findings highlight that the proposed cavity-KC platform offers a concrete and scalable route for exploring the quantum control of MZMs using light. In particular, braiding-enabled quantum teleportation protocols become feasible in this architecture, with immediate relevance to current technologies in superconducting circuits and microwave resonators. This positions our model as a practical framework for testing non-Abelian statistics of free Majorana fermions in nanowires or engineered quantum dot arrays.
From an experimental perspective, recent developments in superconducting quantum circuits and quantum dot systems [6, 66] strongly support the feasibility of realizing the proposed setup. KC-like behavior has already been demonstrated in chains of coupled quantum dots, with parameter control sufficient to access the poor man’s Majorana regime. Coupling these chains to transmission line resonators and applying tailored drives on either the cavity mode or KC sites can allow detection of the predicted phenomena via the transmission spectrum [27, 28, 59, 60], quantum process tomography [61], or microwave interferometry [62].
Finally, we foresee multiple directions for further research. An intriguing avenue is the generalization to multi-cavity architectures, where geometric phase-based braiding mechanisms could be controlled with high precision. Additionally, exploring the effects of photon losses, dephasing, and other sources of environmental dissipation will be crucial for assessing the robustness of light-induced Majorana processes and for guiding their realization in realistic experimental setups.
Acknowledgments
@@
L.Q. and F.J.R. would like to thank Facultad de Ciencias-UniAndes projects: INV-2021-128-2292 and INV-2023-162-2833 for financial support. F.J.G-R and I.A.B-G acknowledge financial support from Spanish MCIN with funding from European Union Next Generation EU (PRTRC17.I1) and Consejeria de Educacion from Junta de Castilla y Leon through QCAYLE project. Additional support from the Department of Education, Junta de Castilla y Leon, and FEDER funds (CLU-2023-1-05) is also acknowledged. C.T. acknowledges financial support from the project “Nanofotonica para Computacion Cuantica” (NanoQuCo) with reference Y2020/TCS-6545 of the Comunidad de Madrid (CAM).
Appendix A Extended properties of two bulk sites case
Similarly to the single site cavity case, in the weak-coupling regime matter states given by Eq. (8) are suited best for exploring the dynamics controlled by $\hat{H}_{2}$. Nevertheless, in the USC regime a better-adapted basis set is provided by occupation number states, $\{|n_{L,R},n_{c1},n_{c2}\rangle\}$, of a non-local fermion $f_{L,R}=\frac{1}{2}\left(\hat{\gamma}_{L}+i\hat{\gamma}_{R}\right)$ and two local fermions fully contained inside the cavity, $c_{1}=\frac{1}{2}\left(\hat{\gamma}_{s,1}+i\hat{\gamma}_{s,2}\right)$ and $c_{2}=\frac{1}{2}\left(\hat{\gamma}_{s+1,1}+i\hat{\gamma}_{s+1,2}\right)$. Now the basis transformation is:
$|-,\mp\rangle_{z}$ $=\frac{1}{2}\Bigl{[}|\circ_{L,R}\rangle\otimes|\bar{\Psi}_{1}^{\pm}\rangle-i|\bullet_{L,R}\rangle\otimes|\bar{\Psi}_{2}^{\pm}\rangle\Bigl{]},$
$|+,\pm\rangle_{z}$ $=\frac{1}{2}\Bigl{[}|\circ_{L,R}\rangle\otimes|\bar{\Psi}_{1}^{\pm}\rangle+i|\bullet_{L,R}\rangle\otimes|\bar{\Psi}_{2}^{\pm}\rangle\Bigl{]}.$ (16)
where $|\bar{\Psi}_{1}^{\pm}\rangle=|\circ_{c_{1}}\circ_{c_{2}}\rangle\pm|\bullet_{c_{1}},\bullet_{c_{2}}\rangle$ and $|\bar{\Psi}_{2}^{\pm}\rangle=|\circ_{c_{1}},\bullet_{c_{2}}\rangle\pm|\bullet_{c_{1}},\circ_{c_{2}}\rangle$. In the USC regime, the ground state is simply expressed as $|GS\rangle\simeq|n_{L,R}\rangle\otimes|\Psi_{c_{1},c_{2},ph}\rangle$ where $|\Psi_{c_{1},c_{2},ph}\rangle$ is a composite state of fermions inside the cavity $c_{1}$ and $c_{2}$, as well as cavity photons. Different regimes of relative values of $\lambda_{1}$ and $\lambda_{2}$ should be discriminated. Given $\lambda_{1}>0$, three different relative coupling sectors should be separately considered, as:
$|{\rm GS}\rangle$ $=|\circ_{L,R},\bullet_{c_{1}},\bullet_{c_{2}}\rangle\otimes|\alpha_{\rm I}\rangle,\;\quad{\rm for}\;\lambda_{2}>0,$ (17)
$|{\rm GS}\rangle$ $=i|\bullet_{L,R},\circ_{c_{1}},\bullet_{c_{2}}\rangle\otimes|\alpha_{\rm II}\rangle,\;\;{\rm for}\;-\lambda_{1}<\lambda_{2}<0,$ (18)
$|{\rm GS}\rangle$ $=i|\bullet_{L,R},\bullet_{c_{1}},\circ_{c_{2}}\rangle\otimes|\alpha_{\rm III}\rangle,\;{\rm for}\;\lambda_{2}<-\lambda_{1}.$ (19)
Here, the coherent state displacements are given by $\alpha_{\rm I}=-(\lambda_{1}+\lambda_{2})/(\sqrt{2}\omega)$, $\alpha_{\rm II}=-\lambda_{1}/(\sqrt{2}\omega)$, and $\alpha_{\rm III}=-\lambda_{2}/(\sqrt{2}\omega)$.
Figure 6: Ground state indicators for the two bulk sites case. In the left panel, the symbols represent the ground state energy obtained by exact diagonalization of the Hamiltonian given by Eq. (9) with $\lambda_{1}=4$ for the resonance case. The solid lines represent the USC approximation given by $E_{\rm GS}=-\omega|\alpha|^{2}$, using $\alpha_{\rm III}$ [Eq. (19)] in the blue region, $\alpha_{\rm II}$ [Eq. (18)] in the red region, and $\alpha_{\rm I}$ [Eq. (18)] in the green region. The dashed black line corresponds to the exact results by first-order perturbation given by Eq. (20). In the right panel, using the same parameters as the left panel, we contrast the numerical results (symbols) and the exact results by employing first-order perturbation theory (dashed lines) for the ground state fidelity. The blue/red symbols correspond to the numerical results obtained using Eq. (17) and Eq. (18), respectively. For the exact first-order perturbation results, we used Eq. (21), with the corresponding angle given by Eq. (22).
Numerical results for the ground state energy in each sector are depicted in Figs. 6-7, showing a good agreement with predicted USC analytical results, $E_{GS}=-|\alpha|^{2}\omega$. A crossing behavior is evident at $\lambda_{2}=-\lambda_{1}$ due to the quasi-orthogonality of coherent states $|\alpha\rangle$ and $|-\alpha\rangle$ for large values of $\lambda$, whereas an anti-crossing behavior takes place at $\lambda_{2}=0$. In the latter case, convenient approximate expressions can be found for both energies and ground state composition by considering the free fermion term in the Hamiltonian $\hat{H}_{2}$, Eq. (9), as a perturbation acting on the sub-space generated by states $\{|GS_{+}\rangle,GS_{-}\rangle\}$. The ground state energy up to the first order in $\Delta$ is given by:
$E_{\rm GS}=-\frac{1}{2}\Bigl{[}\omega\left(\alpha_{\rm I}^{2}+\alpha_{\rm II}^{2}\right)+\sqrt{\omega^{2}\left(\alpha_{\rm I}^{2}-\alpha_{\rm II}^{2}\right)^{2}+4\Delta^{2}|\langle\alpha_{\rm II}|\alpha_{\rm I}\rangle|^{2}}\Bigr{]},$ (20)
while the ground state in the transition regime becomes expressed as the superposition:
$|{\rm GS}\rangle=-\sin\Bigl{(}\frac{\theta}{2}\Bigr{)}|{\rm GS}\rangle_{\rm I}+\cos\Bigl{(}\frac{\theta}{2}\Bigr{)}|{\rm GS}\rangle_{\rm II},$ (21)
where
$\tan\left(\theta\right)=-\frac{4\omega\Delta}{\lambda_{2}\left(\lambda_{2}+2\lambda_{1}\right)}e^{-\frac{\lambda_{2}^{2}}{4\omega^{2}}}.$ (22)
Thus, it is evident that varying the coupling strength $\lambda_{2}$ of one physical site with the cavity and crossing the point $\lambda_{2}=0$, a cavity allowed flip-flop effect between the topological fermion state $|n_{L,R}\rangle$ and the local fermion $|n_{c1}\rangle$ is attained.
Figure 7: Parameter Regimes for the Ground State Structure and Braiding in a 2CK Setup. In the left panel, we present a schematic representation of the ground state structure for 2KC as a function of $\lambda_{1}$ and $\lambda_{2}$ [see Eqs. (17), (18), and  (19)]. Additionally, we validate the theoretical results for the ground state structure by numerically evaluating the expectation values of the operators $\hat{a}^{\dagger}\hat{a}$ and $\hat{a}$, and contrasting these numerical results with the theoretical predictions. In the right panel, we illustrate the landscape of parity between $\gamma_{L}$ and $\gamma_{R}$. The solid blue line corresponds to the case where $\lambda_{2}$ is real and changes from positive to negative regimes. The closed yellow ring shows the evolution of the parity, demonstrating the perfect braiding between the Majorana fermions $\gamma_{L}$ and $\gamma_{R}$, but now, in contrast to the main text, we used both outside MFs.
Appendix B Braiding protocol in a two-site cavity
We examine the possibility of reaching the same $e^{-i\frac{\pi}{4}}$ on the GS by performing an adiabatic cyclic evolution of the coherent cavity field (Berry phase). The ground state (GS) adiabatic evolution is fully contained in the two-dimensional sub-space spanned by the states $\{|{\rm GS}_{+}\rangle,|{\rm GS}_{-}\rangle\}$, as previously assured. Thus, the effective Hamiltonian matrix in that sub-space becomes:
$\hat{H}_{2}=\left(\begin{array}[]{cc}-\omega|\alpha_{+}|^{2}&\Delta\langle\alpha_{-}|\alpha_{+}\rangle\\ \Delta\langle\alpha_{+}|\alpha_{-}\rangle&-\omega|\alpha_{-}|^{2}\end{array}\right).$ (23)
With $\alpha_{+}=-\frac{\lambda_{1}+\lambda_{2}e^{i\phi_{2}}}{\sqrt{2}\omega}$ and $\alpha_{-}=-\frac{\lambda_{1}}{\sqrt{2}\omega}$. We perform a cyclic evolution of the GS by fixing a $\lambda_{2}>0$ value while varying the phase $\phi_{2}=\phi$ of the coupling strength, $0\leq\phi\leq 2\pi$. Thus, for the GS we have
$|{\rm GS}(\phi)\rangle=C_{+}(\phi)|{\rm GS}_{+}(\phi)\rangle+C_{-}(\phi)|{\rm GS}_{-}\rangle.$ (24)
Notice that the state $|{\rm GS}_{-}\rangle$ does not depend on the phase $\phi$. We proceed with an analytical calculation of the Berry phase using Eq.(23) and Eq.(24). The Berry phase is given by:
$\varphi_{B}=i\int_{0}^{2\pi}d\phi\langle{\rm GS}(\phi)|\frac{\partial}{\partial\phi}|{\rm GS}(\phi)\rangle.$ (25)
Inserting Eq. (24) into Eq. (25) one gets Eq. (13).